# Conduct Parameter Estimation in Homogeneous Goods Markets with Equilibrium Existence and Uniqueness Conditions: The Case of Log-linear Specification

### v1

@@ Abstract
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in a log-linear model with homogeneous goods markets. First, we derive such conditions. Second, Monte Carlo simulations confirm that in a log-linear model, incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters.
Keywords: Conduct parameters, Homogenous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
@@ 1 Introduction
Measuring competitiveness is a crucial task in the empirical industrial organization literature. Conduct parameter is considered a useful measure of competitiveness. However, it cannot be directly measured from data because data usually lack information about marginal cost. Therefore, researchers aim to identify and estimate the conduct parameter.
As the simplest specification, Bresnahan (1982) considers identification of conduct parameters for the linear model. Matsumura and Otani (2023b, a) resolves the conflict on some identification problems between Bresnahan (1982) and Perloff and Shen (2012). However, researchers often implement alternative specifications, such as the log-linear model (Okazaki et al. 2022, Mérel 2009). In the context of the log-linear model, these papers identify estimation issues, with some estimated conduct parameters being unrealistically low or even negative. This raises doubts about the methodology, despite the identification strategy outlined by Lau (1982). This is an obstacle to choosing a better specification of the demand and supply functions.
To overcome the problem, we propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices as constraints. First, we prove that a unique equilibrium exists under certain theoretical conditions. As far as we know, this is a new result. Second, we show that incorporating equilibrium existence and uniqueness conditions resolves the above estimation problems and makes the model work in estimating conduct parameters even in the log-linear model. Hence, our results support those of Lau (1982) numerically.
@@ 2 Model
$P_{t}=-\theta\frac{\partial P_{t}(Q_{t})}{\partial Q_{t}}Q_{t}+MC_{t}(Q_{t}),$ (1)
where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$, which is the conduct parameter. The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ 1)1)1 ) (Bresnahan 1982).
Consider an econometric model. Assume that the demand and the marginal cost functions are written as follows:
@@
where $X^{d}_{t}$ and $X^{c}_{t}$ are the vector of exogenous variables, $\varepsilon^{d}_{t}$ and $\varepsilon^{c}_{t}$ are the error terms, and $\alpha$ and $\gamma$ are the vector of parameters. We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
@@ 2.1 Log-linear demand and log-linear marginal cost
Consider a log-linear model, which is a typical specification. The demand and marginal cost functions are specified as
@@
where $Y_{t}$ is an excluded demand shifter, $W_{t}$ and $R_{t}$ are excluded cost shifters, and $Z_{t}^{R}$ is Bresnahan’s demand rotation instrument. Then, Equation (1) is written as
@@
By taking logarithm of Equation (4) and substituting Equation (3), we obtain
@@
In this model, the number of equilibrium prices varies. Although it is a widely known model, to our knowledge, there is no paper examining this. Thus, we derive the conditions for the uniqueness in the following proposition which are used in our estimation.
@@
Assume that $\alpha_{1}+\alpha_{2}Z^{R}_{t}\neq 0$. Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$. The number of equilibrium prices $P_{t}^{*}>0$ is determined as follows:
  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, there is no equilibrium price,
  • When $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\neq 1$, there is a unique equilibrium price,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, there are infinitely many equilibrium prices when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium price otherwise.
See the online appendix A.1 for the proof.
@@ 3 Estimation
To estimate parameters in the demand and supply equations, we use GMM estimation. Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using Equation (2) and (5).
Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of parameters. Given the demand equation (2) and the supply equation (5), we can write the error terms in the demand and supply equation as
@@
To estimate the parameters, we convert the conditional moment conditions, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moment conditions, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$. Using Equations (6) and (7), we construct the sample analog of the moment conditions:
@@
We define the GMM estimator as the vector $\xi^{*}$ that solves the problem,
$\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
@@
The solution to (8) is called the N2SLS estimator (Wooldridge 2010).
Using Proposition 1, we impose the following constraints:
@@
Constraint (12) is a standard assumption on the conduct parameter. Constraints (13) and (14) relate to the uniqueness of equilibrium prices. Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost. See online appendix A.2 for the detailed setting.
@@ 4 Simulation results
We compare N2SLS estimation with and without Constraints (12), (13), and (14). Table 1 presents the results. First, as reported in the literature (Mérel 2009, Okazaki et al. 2022), N2SLS without constraints converges to implausibly low and negative conduct parameters. This is because the algorithm searches the area in which no equilibrium price and quantity exist and then finds unreasonable local optima that might have a lower objective function than at the true values. Second, N2SLS with constraints converges to the reasonable values. As the sample size increases, the bias and root-mean-square error (RMSE) decrease to levels comparable to linear models. Third, we find an ad hoc method that uses Equation (3) to compute the residual in the supply estimation, $\varepsilon_{t}^{c}$, and Equation (4) as an equality constraint with Constraints (12), (13), and (14). Then, the bias and RMSE of conduct parameter $\theta$ are reduced to 0.014 and 0.217 in Panel (c), although the results do not dominate the results in Panel (b) for all parameters. Therefore, incorporating equilibrium existence and uniqueness conditions simply is helpful for conduct parameter estimation in nonlinear models such as a log-linear model. See online appendix A.4 for additional experiments under different variances of errors and the results of a linear model.
@@
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030
$\alpha_{1}$0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574
$\alpha_{2}$0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043
$\alpha_{3}$0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187
$\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351
$\gamma_{1}$0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220
$\gamma_{2}$0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127
$\gamma_{3}$0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133
$\theta$3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05
Runs converged (%) 99.500 99.800 98.600 98.400
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563
$\alpha_{1}$0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241
$\alpha_{2}$0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019
$\alpha_{3}$0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152
$\gamma_{0}$1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570
$\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170
$\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116
$\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118
$\theta$0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333
Runs converged (%) 99.600 99.900 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
@@
Note: The error terms are drawn from a normal distribution, $N(0,\sigma)$. True values: $\alpha_{0}=20.0,\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=1.0,\gamma_{0}=5.0,\gamma_{1}=1.0,\gamma_{2}=1.0,\gamma_{3}=1.0,\theta=0.5$ and $\sigma=1.0$. See online appendix A.2 for the setting.
5 Conclusion
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in homogeneous goods markets. First, we derive the conditions. Second, Monte Carlo simulations confirm that incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters in a log-linear model, as reported in the literature.

### v2

@@ Abstract
We propose a constrained generalized method of moments (GMM) estimator with some equilibrium uniqueness conditions for estimating the conduct parameter in a log-linear model with homogeneous goods markets. Monte Carlo simulations demonstrate that merely imposing parameter restrictions leads to not just inaccurate estimations but also some numerical issues, and adding the equilibrium uniqueness conditions resolves them. We also suggest a formulation of the GMM estimation to further avoid the numerical issues.
Keywords: Conduct parameters, Homogeneous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
@@ 1 Introduction
Measuring competitiveness is central in empirical IO, and the conduct parameter is a widely used proxy. Since marginal cost is typically unobserved, researchers identify and estimate the conduct parameter indirectly.
bresnahan1982oligopoly develops an identification strategy under linear demand and marginal cost, and matsumura2023resolving provide detailed conditions for this case. Yet, empirical work often employs log-linear models (okazaki2022excess, merel2009measuring), which tend to produce implausibly low or negative estimates, even though identification conditions for general models are available from lau1982identifying and matsumura2025revisiting. This casts doubt on the methodology and complicates model selection.
We address this by proposing a constrained GMM estimator that incorporates theoretical conditions for the uniqueness of equilibrium. First, we derive new conditions guaranteeing a unique equilibrium. Second, Monte Carlo simulations show that parameter restrictions alone yield inaccurate estimates and numerical errors, while adding equilibrium conditions resolves them. We also propose a modified GMM formulation that further mitigates these issues.
@@ 2 Model
$P_{t}+\theta Q_{t}P_{t}^{\prime}(Q_{t})=MC_{t}(Q_{t}),$ (1)
where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the inverse demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$ is the conduct parameter. The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ $1)$. See bresnahan1982oligopoly for the details.
Consider an econometric model. Assume that the inverse demand and the marginal cost functions are given as
@@
where $X^{d}_{t}$ and $X^{c}_{t}$ are the vector of exogenous variables, $\varepsilon^{d}_{t}$ and $\varepsilon^{c}_{t}$ are the error terms, and $\alpha$ and $\gamma$ are the vector of parameters. We allow $X^{d}_{t}$ and $X^{c}_{t}$ to have common variables, but assume that there is at least one demand variable and one cost variable that are mutually excluded. We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition, $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
The identification of the conduct parameter is indirectly characterized by lau1982identifying:
Theorem 1.
Under the assumption that the industry inverse demand and cost functions are twice continuously differentiable, the index of competitiveness $\theta$ cannot be identified from data on industry price and output and other exogenous variables alone if and only if the industry inverse demand function is separable in $X^{d}$, that is, $f(Q,r(X^{d}))$, but not take the form $P=Q^{-1/\theta}r(X^{d})+s(Q)$.
This theorem implies that the conduct parameter is identified if the inverse demand function is not separable. A demand rotation instrument (bresnahan1982oligopoly) achieves this. See Appendix A.1 for the details of the definition of separability.
@@ 2.1 Log-linear demand and log-linear marginal cost
Consider a log-linear model, which is a typical specification. The inverse demand and marginal cost functions are specified as
@@
where $Y_{t}$ and $Z_{t}^{R}$ are excluded demand shifters and $W_{t}$ and $R_{t}$ are excluded cost shifters. When $Y_{t}$ and $Z_{t}^{R}$ vary without changing the equilibrium quantity, they work as the demand rotation instrument. Then, (1) is written as
@@
By taking logarithm of (4) and substituting (3), we obtain
@@
The intersection of (2) and (5) determines the equilibrium, but there could be multiple equilibria. Although this model is widely known, no paper has examined the multiple equilibria problem to our knowledge. The next proposition provides the conditions for uniqueness. The proof is in the online appendix A.2.
@@
Assume that $\alpha_{1}+\alpha_{2}Z^{R}\neq 0$. Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y+\varepsilon^{d}}{\alpha_{1}+\alpha_{2}Z^{R}}+\gamma_{2}\log W+\gamma_{3}\log R+\varepsilon^{c}$. The number of equilibria is determined as follows:
  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})\leq 0$, there is no equilibrium,
  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})>0$,
    • If $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$, there is a unique equilibrium,
    • If $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}=0$, there are infinitely many equilibria when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium otherwise.
The condition $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})>0$ rules out the region where the log transformation leaves its domain, which corresponds to implausibly elastic demand combined with large conduct. The assumption $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$ excludes the knife-edge case in which the (pseudo) supply and demand are exactly parallel so that every price could be an equilibrium; both restrictions are technical and do not bind in regular empirical settings.
@@ 3 Estimation
Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of the parameters in the model. We use the GMM for the estimation. Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using (2) and (5). We rewrite the demand equation (2) and the supply equation (5) as
@@
To estimate the parameters, we convert the conditional moments, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moments, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$. Using Equations (6) and (7), we construct the sample analog of the unconditional moments:
@@
We define the N2SLS estimator as the solution to the problem,
$\xi^{*}=\arg\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
@@
We also add the following constraints based on Proposition 1 to (8):
@@
Constraint (12) is a standard assumption on the conduct parameter. Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost, which guarantees that $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$. Constraint (14) relates to the uniqueness of equilibrium. See the detailed simulation setting in the online appendix A.3.
@@ 4 Simulation results
We compare N2SLS estimations with and without constraints in Table 1. Panel (a) shows that, without constraints, the estimator fails to recover $\gamma_{0}$ and $\theta$, replicating known issues due to the flat objective function and invalid search regions without equilibrium (Appendix A.4). Panel (b), which imposes Constraint (12), improves estimation in large samples via the domain restriction. [^1] Constraints (13) and (14) alone yield severe bias; see Table 9 and Appendix A.5. However, in small samples, demand parameter estimates degrade and convergence declines. When convergence fails, $\alpha_{1}$ becomes large, rendering $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})<0$ and causing numerical errors inside the log term in (1). Adding constraints (13) and (14) in Panel (c) improves small-sample convergence and demand accuracy, though convergence is not guaranteed. In large samples, performance surpasses that of Panel (b) for some parameters.
To address convergence failure, we propose an alternative formulation (Table 2) that computes $\varepsilon^{c}_{t}$ via (3) and enforces Equation (4) as a constraint, along with Constraints (12)–(14). [^2] See Appendix A.3 for details. This avoids log terms in both objective and constraints, achieving 100% convergence and reducing $\theta$’s bias and RMSE to 0.014 and 0.217, though not dominating Panel (c) across all parameters. In sum, incorporating equilibrium uniqueness conditions and eliminating log terms greatly improves conduct parameter estimation. Additional experiments appear in Appendix A.5.
@@
(a) N2SLS without Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030 $\alpha_{1}$ -0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574 $\alpha_{2}$ -0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043 $\alpha_{3}$ -0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187 $\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351 $\gamma_{1}$ -0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220 $\gamma_{2}$ -0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127 $\gamma_{3}$ -0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133 $\theta$ -3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05 Runs converged (%) 99.500 99.800 98.600 98.400 Sample size ($T$) 100 200 1000 1500
(b) N2SLS with Constraints (12)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.922 8.603 -0.068 5.116 0.037 2.035 0.000 1.556 $\alpha_{1}$ -0.299 1.314 -0.010 0.785 0.005 0.312 0.000 0.240 $\alpha_{2}$ -0.013 0.104 -0.002 0.063 0.001 0.024 0.000 0.019 $\alpha_{3}$ -0.165 0.774 -0.007 0.472 -0.004 0.185 -0.001 0.152 $\gamma_{0}$ -1.767 14.394 -1.001 6.530 -0.208 1.993 -0.156 1.566 $\gamma_{1}$ 0.255 1.949 0.132 0.838 0.034 0.229 0.027 0.174 $\gamma_{2}$ 0.125 1.097 0.053 0.475 0.017 0.150 0.019 0.119 $\gamma_{3}$ 0.099 0.903 0.062 0.481 0.007 0.149 0.014 0.120 $\theta$ -0.098 0.441 -0.060 0.421 -0.061 0.319 -0.058 0.281 Runs converged (%) 98.100 98.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
(c) N2SLS with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563 $\alpha_{1}$ -0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241 $\alpha_{2}$ -0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019 $\alpha_{3}$ -0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152 $\gamma_{0}$ -1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570 $\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170 $\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116 $\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118 $\theta$ -0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333 Runs converged (%) 99.600 99.900 100.000 100.000 Sample size ($T$) 100 200 1000 1500
@@
Note: The error terms are drawn from a normal distribution, $N(0,\sigma)$. True values: $\alpha_{0}=20.0,\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=1.0,\gamma_{0}=5.0,\gamma_{1}=1.0,\gamma_{2}=1.0,\gamma_{3}=1.0,\theta=0.5$ and $\sigma=1.0$. See online appendix A.3 and matsumura2024challenges for the setting.
Table 2: Ad hoc method using (3) to compute $\varepsilon_{t}^{c}$ and (4) with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555 $\alpha_{1}$ -0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240 $\alpha_{2}$ -0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020 $\alpha_{3}$ -0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152 $\gamma_{0}$ -5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533 $\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175 $\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120 $\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121 $\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217 Runs converged (%) 100.000 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
5 Discussion
Two concerns surround the conduct parameter approach: the difficulty of interpreting intermediate or extreme values, and the critique by corts1999conduct that it may understate market power under collusion. [^3] As magnolfi2022comparison note, this critique does not apply when the data stem from a static model. We show that implausible estimates in log-linear models often stem from numerical issues—especially when equilibrium conditions are omitted—rather than conceptual flaws. Addressing these issues yields more stable and interpretable results.
This distinction matters: misattributing numerical artifacts to theoretical limits risks dismissing a useful tool. Our findings aim to encourage more constructive use of the conduct parameter approach.